# SpecExec: Massively Parallel Speculative Decoding for Interactive LLM Inference on Consumer Devices

### v1

@@ 2.2 Parameter Offloading
Another recent line of work explores running and training large models with limited accelerator memory by “offloading” their parameters to another, cheaper memory, such as RAM or even SSD storage [16, 17, 18]. This technique works by loading model parameters on the GPU when they are needed for computation. Since most deep learning models use layers in a fixed order, offloading can pre-dispatch the next layer parameters in the background ahead of time.
This technique works particularly well when processing large batches of data, during training [16, 17] or large-batch non-interactive inference [19, 20], where each layer process multiple tokens each time it is loaded. In turn, when doing interactive inference, offloading works significantly slower than on-device inference. This is because interactive inference has to process one or few tokens at a time, and therefore spends most of the time waiting for the parameters to load.

### v2

@@ 2.2 Parameter Offloading
Offloading has emerged as a crucial technique for leveraging GPUs with insufficient on-chip memory to hold entire models, to perform training or inference with these models. When using offloading, the model parameters are typically stored in CPU RAM, and layers are sequentially loaded onto the GPU when they are needed during a forward or backward pass  [16, 17, 18]. This method is most effective when the model is able to process many tokens in parallel (e.g., during training or large-batch inference), where the overhead of transferring layers between CPU and GPU is amortized over many computations  [19, 20]. However, it is less effective for small-batch inference, because generating each new token requires transferring the entire model from CPU to GPU, which is heavily bottlenecked by the memory bandwidth (even with PCIe 4.0).